# Packing Input Frame Context in Next-Frame Prediction Models for Video Generation

### v2

@@ Abstract
We present a neural network structure, FramePack, to train next-frame (or next-frame-section) prediction models for video generation. The FramePack compresses input frames to make the transformer context length a fixed number regardless of the video length. As a result, we are able to process a large number of frames using video diffusion with computation bottleneck similar to image diffusion. This also makes the training video batch sizes significantly higher (batch sizes become comparable to image diffusion training). We also propose an anti-drifting sampling method that generates frames in inverted temporal order with early-established endpoints to avoid exposure bias (error accumulation over iterations). Finally, we show that existing video diffusion models can be finetuned with FramePack, and their visual quality may be improved because the next-frame prediction supports more balanced diffusion schedulers with less extreme flow shift timesteps.
@@ 1 Introduction
Forgetting and drifting are the two most critical problems in next-frame (or next-frame-section) prediction models for video generation. Herein, "forgetting" refers to the fading of memory as the model struggles to remember earlier content and maintain consistent temporal dependencies, whereas "drifting" refers to the iterative degradation of visual quality due to error accumulation over time (also called exposure bias).
@@
A fundamental dilemma emerges when attempting to simultaneously address both forgetting and drifting: any method that mitigates forgetting by enhancing memory may also make error accumulation/propagation faster, thereby exacerbating drifting; any method that reduces drifting by interrupting error propagation and weakening the temporal dependencies (e.g., masking or re-noising the history) may also worsen the forgetting. This essential trade-off hinders the scalability of next-frame prediction models.
@@
The forgetting problem leads to a naive solution to encode more frames, but this quickly becomes computationally intractable due to the quadratic attention complexity of transformers (or sub-quadratic optimizations like FlashAttn, etc.). Moreover, video frames contain significant temporal redundancy, rendering naive full-context approaches less efficient. The substantial duplication of visual features across consecutive frames reveals the potential to design effective compression systems to facilitate memorization.
@@
The drifting problem is influenced by memorizing mechanisms from multiple aspects. The source of drifting lies in initial errors that occur in individual frames, while the effect is the propagation and accumulation of these errors across subsequent frames, leading to degraded visual quality. On one hand, a stronger memorizing mechanism leads to better temporal consistency and reduces the occurrence of initial errors, thereby mitigating drifting. On the other hand, a stronger memorizing mechanism also memorizes more errors and thus accelerates error propagation when errors do occur, exacerbating drifting. This paradoxical relationship between memory mechanisms and drifting necessitates carefully designed training/sampling methods to facilitate error correction or interrupt error propagation.
@@
In this paper, we propose FramePack as an anti-forgetting memory structure along with anti-drifting sampling methods. The FramePack structure addresses the forgetting problem by compressing input frames based on their relative importance, ensuring the total transformer context length converges to a fixed upper bound regardless of video duration. This enables the model to encode significantly more frames without increasing the computational bottleneck, thereby facilitating anti-forgetting. Moreover, we propose anti-drifting sampling methods that break the causal prediction chain and incorporate bi-directional context. These methods include generating endpoint frames before filling intermediate content, and an inverted temporal sampling approach where frames are generated in reverse order, each attempting to approach a known high-quality frame (e.g., an input frame in image-to-video tasks). We show that these methods effectively reduce the occurrence of errors and prevent their propagation.
@@
We demonstrate that existing pretrained video diffusion models (e.g., HunyuanVideo, Wan, etc.) can be finetuned with FramePack. Our experiments reveal several findings: because next-frame prediction generates smaller tensor sizes per step compared to full-video generation, it enables more balanced diffusion schedulers with less extreme flow shift timesteps. We observe that the resulting less aggressive schedulers may lead to improved visual quality, beyond the direct goals of addressing the forgetting and drifting problems.
@@ 2.1 Anti-forgetting and Anti-drifting
Noise scheduling and augmentation in history frames modify noise levels at specific timesteps, video times, or image frequencies to create causal computation or anti-drifting effects. These methods generally reduce dependency on past frames. DiffusionForcing [4] and RollingDiffusion [32] are typical examples. Our ablation studies investigate the influence of adding noise to history frames.
@@
Compressing latent space can improve the efficiency of video diffusion models. LTXVideo [13] shows that a highly compressed latent space can be used for diffusing videos efficiently. PyramidFlow [20] diffuses video latents in a pyramid and renoise downsampled latents in that pyramid to reduce computation costs. FAR [12] proposes a multi-level causal attention structure to establish long-short-term causal context pacifying and KV caches. HiTVideo [62] uses hierarchical tokenizers to enhance the video generation with autoregressive language models.
The trade-off between forgetting and drifting is also evidenced by previous discussions. CausVid [53] shows that when the video generator is causal, the quality degradation appears at the end of the video and the final video length may be subject to an upper bound. DiffusionForcing [4] discussed that the cause of this drift may be related to error accumulation in models’ observation disparity between training and inference. Wang et al. [40] discussed that a model with stronger memory may suffer more from drifting and error accumulation.
@@ 2.2 Long Video Generation
Extending video generation beyond short clips remains an open problem. LVDM [14] generates long videos using latent diffusion, while Phenaki [38] creates variable-length videos from sequences of text prompts. Gen-L-Video [39] applies temporal co-denoising for multi-text conditioned videos, and FreeNoise [31] extends pretrained models without additional training via noise rescheduling. NUWA-XL [50] implements a Diffusion-over-Diffusion architecture with coarse-to-fine processing, while Video-Infinity [36] overcomes computational constraints through distributed generation. StreamingT2V [15] produces consistent, dynamic, and extendable videos without hard cuts, and CausVid [53] transforms bidirectional models into fast autoregressive ones through distillation. Recent advances include GPT-like architecture (ViD-GPT [11]), multi-event generation (MEVG [30]), attention control for multi-prompt generation (DiTCtrl [3]), precise temporal control (MinT [44]), history-based guidance (HistoryGuidance [34]), unified next-token and full-sequence diffusion (DiffusionForcing [4]), SpectralBlend temporal attention (FreeLong [27]), video autoregressive modeling (FAR [12]), and test-time training (TTT [6]).
2.3 Efficient Architectures for Video Generation
We discuss typical methods to improve video model efficiency. Linear attention [2, 47, 41, 5, 54, 21] reduces the attention complexity by reformulating linear operations. Sparse attention [45, 57, 58, 46] computes attention only on important token pairs, skipping negligible attention values. Low-bit computation [24, 60, 23] quantizes model weights and activations to lower precision for faster computation. Low-bit attention [56, 55] specifically optimizes attention computation through quantization techniques. Hidden state caching [29, 25] reuses intermediate computations across diffusion timesteps to avoid redundant calculations. Distillation [35, 28, 51, 52] transfers knowledge from larger models to smaller ones or reduces sampling steps while preserving quality.
3 Method
@@
We consider a video generation model that predicts next frames repeatedly to form a video. For simplicity, we consider diffusion-based next-frame-section prediction models using Diffusion Transformers (DiTs) that generate a section of $S$ unknown frames $\bm{X}\in\mathbb{R}^{S\times h\times w\times c}$ conditioned on $T$ input frames $\bm{F}\in\mathbb{R}^{T\times h\times w\times c}$. All definitions of frames and pixels refer to latent representations, as most modern models operate in latent space.
@@
For next-frame (or next-frame-section) prediction, $S$ is typically 1 (or a small number). We focus on the challenging case where $T\gg S$. With per-frame context length $L_{f}$ (typically $L_{f}\approx 1560$ for each 480p frame in Hunyuan/Wan/Flux), the vanilla DiT yields total context length $L=L_{f}(T+S)$. This causes context length explosion when $T$ is large.
3.1 FramePack
@@
We observe that the input frames have different importance when predicting the next frame, and we can prioritize the input frames according to their importance. Without loss of generality, we consider a simple case where the temporal proximity reflects importance: frames temporally closer to the prediction target can be more relevant. We enumerate all frames with $\bm{F}_{0}$ being the most important (e.g., the most recent) and $\bm{F}_{T-1}$ being the least (e.g., the oldest).
We define a length function $\phi(\bm{F}_{i})$ that determines each frame’s context length after VAE encoding and transformer patchifying, applying progressive compression to less important frames as
@@
where $\lambda>1$ is a compression parameter. The frame-wise compression is achieved by manipulating the transformer’s patchify kernel size in the input layer (e.g., $\lambda=2,i=5$ means a kernel size where the product of all dims equals $322^{5}=$ 32 like the 3D kernel $2\times 4\times 4$, or $8\times 2\times 2$, etc.). The total context length then follows a geometric progression
@@
and when $T\to\infty$, the total context length converges to
$\lim_{T\to\infty}L=S\cdot L_{f}+\frac{1}{1-1/\lambda}\cdot L_{f}=\left(S+\frac{\lambda}{\lambda-1}\right)\cdot L_{f}\,,$ (3)
and this bounded context length makes FramePack’s computation bottleneck invariant to the input frame number $T$.
Since most hardware supports efficient matrix processing by powers of 2, we mainly discuss the case of $\lambda=2$ in this paper. Note that we can represent arbitrary compression rates by duplicating (or dropping) several specific terms in the power-of-2 sequence: considering the accumulation $\sum_{i=0}^{+\infty}\frac{1}{2^{i}}=\frac{2}{2-1}=2$, if we want to loosen it a bit, for example to $2.625$, we can duplicate the terms $\frac{1}{2}$ and $\frac{1}{8}$ so that $\frac{1}{1}+\frac{1}{2}+(\frac{1}{2})+\frac{1}{4}+\frac{1}{8}+(\frac{1}{8})+\frac{1}{16}+...=\frac{1}{2}+\frac{1}{8}+\sum_{i=0}^{+\infty}\frac{1}{2^{i}}=2.625$. Following this, one can cover arbitrary rates by converting the rate value to binary bits and then translate every bit.
The patchifying operations in most DiTs are 3D, and we denote the 3D kernel as $(p_{f},p_{h},p_{w})$ representing the steps in frame number, height, and width. A same compression rate can be achieved by multiple possible kernel sizes, e.g., the compression rate of $64$ can be achieved by $(1,8,8)$, $(4,4,4)$, $(16,2,2)$, $(64,1,1)$, etc. These would lead to different FramePack compression schedules.
Figure 1: Ablation Variants of FramePack. We present several typical kernel structures of FramePack with commonly used kernel sizes and compression rates. This list does not necessarily cover all popular variants, and more structures can be developed in a similar way.
@@
We observe that features from deep neural networks at different compression rates exhibit notable differences. Empirical evidence shows that using independent parameters for the different input projections at multiple compression rates facilitates stabilized learning. We assign the most commonly used input compression kernels as independent neural network layers: $(2,4,4)$, $(4,8,8)$, and $(8,16,16)$. For higher compressions (e.g., at $(16,32,32)$), we first downsample (e.g., with $2\times 2\times 2$) and then use the largest kernel $(8,16,16)$. This allows us to handle all compression rates. When training these new input projection layers, we initialize their weights by interpolating from the pretrained patchifying projection (e.g., the $(2,4,4)$ projection of HunyuanVideo/Wan).
@@
While in theory FramePack can process videos of arbitrary length with a fixed, invariant context length, practical considerations arise when the input frame length becomes extremely large. In the tail area, frames may fall below a minimum unit size (e.g., a single latent pixel). We discuss 3 options to handle the tail: (1) simply delete the tail; (2) allow each tail frame to increase the context length by a single latent pixel; (3) apply global average pooling to all tail frames and process them with the largest kernel. In our tests, the visual differences between these options are relatively negligible. We note that the tail refers to the least important frames, not always the oldest frames (in some cases, we can assign old frames with higher importance).
@@
When encoding inputs with different compression kernels, the different context lengths require RoPE (Rotary Position Embedding) alignment. RoPE generates complex numbers with real and imaginary parts for each token position across all channels, which we refer to as “phase”. RoPE typically multiplies the phase to neural network features channel-wise. To match the RoPE encoding after compression, we directly downsample (using average pooling) the RoPE phases to match the compression kernels.
3.2 FramePack Variants
The above FramePack structure is discussed with the frame-wise importance defined by simple temporal proximity, and the compression schedule with simple geometric progression. We consider more variants for practical applications and optimal quality.
We visualize typical structures in Fig. 1. First, compression levels can be duplicated and combined with higher compression rates. In Fig. 1-(b), a power-of-4 sequence with each level repeated 3 times allows for same kernel sizes in frame width and height, making the compression more compact. Compression can also be applied in the temporal dimension (e.g., using a power-of-two sequence), as in Fig. 1-(c). This method encodes multiple frames in the same tensor, which is naturally aligned with DiT architectures.
We discuss alternative frame-wise importance modeling beyond simple temporal proximity in Fig. 1-(d,e). For instance, we can assign full-length context to the oldest frame, or treat both beginning and ending frames as equally important while applying higher compression to middle frames. These structures are particularly effective for applications like image-to-video generation, where user-provided initial frames carry higher importance.
3.3 Anti-drifting Sampling

### v3

# Frame Context Packing and Drift Prevention in Next-Frame-Prediction Video Diffusion Models

@@ Abstract
We present a neural network structure, FramePack, to train next-frame (or next-frame-section) prediction models for video generation. FramePack compresses input frame contexts with frame-wise importance so that more frames can be encoded within a fixed context length, with more important frames having longer contexts. The frame importance can be measured using time proximity, feature similarity, or hybrid metrics. The packing method allows for inference with thousands of frames and training with relatively large batch sizes. We also present drift prevention methods to address observation bias (error accumulation), including early-established endpoints, adjusted sampling orders, and discrete history representation. Ablation studies validate the effectiveness of the anti-drifting methods in both single-directional video streaming and bi-directional video generation. Finally, we show that existing video diffusion models can be finetuned with FramePack, and analyze the differences between different packing schedules.
@@ 1 Introduction
Forgetting and drifting are the two most critical problems in next-frame-prediction models for video generation. “Forgetting” refers to the fading of memory as the model struggles to remember earlier content and maintain consistent temporal dependencies. “Drifting” refers to the degradation of visual quality due to error accumulation over time (also called exposure/observation bias).
@@
A fundamental dilemma emerges when attempting to simultaneously address both forgetting and drifting: any method that mitigates forgetting by enhancing memory may also increase error accumulation/propagation, thereby exacerbating drifting; any method that reduces drifting by interrupting error propagation needs to weaken temporal dependencies (e.g., masking or re-noising the history), thus worsening the forgetting. This is a fundamental trade-off that hinders the scalability of next-frame prediction models.
@@
A naive solution to forgetting is to encode more frames. But this approach quickly becomes computationally intractable due to the quadratic attention complexity of transformers (even with optimizations like Flash Attention [10], etc.). Moreover, video frames contain significant temporal redundancy, making naive full-context approaches very inefficient. The substantial duplication of visual features across consecutive frames suggests the potential to design effective compression systems to facilitate memorization.
@@
Drifting is influenced by memorizing mechanisms in several ways. The source of drifting lies in the initial errors that occur in individual frames, while the effect is the propagation and accumulation of these errors across subsequent frames, and the eventual drifting out of the train distribution. A stronger memorization mechanism, on one hand, can lead to better temporal consistency and reduce the occurrence of initial errors. On the other hand, it also memorizes more errors and thus accelerates error propagation when errors do occur. This paradoxical relationship between memory mechanisms and drifting necessitates carefully designed training/sampling methods to facilitate error correction or interrupt error propagation.
@@
In this paper, we propose FramePack as an anti-forgetting memory structure along with anti-drifting sampling and training methods. The FramePack structure addresses the forgetting problem by compressing input frames based on their relative importance, ensuring that the total transformer context length converges to a fixed upper bound. We consider both time-proximity-based and feature-similarity-based importance measures. Afterwards, we propose anti-drifting sampling methods that break the causal prediction chain and incorporate bi-directional contexts by planning single or multiple endpoint frames. We also present an anti-drifting training method to convert frame history into discrete tokens so as to reduce the history disparity between training and inference. We show that these methods effectively reduce the occurrence of errors and prevent their propagation.
@@
We demonstrate that existing pretrained video diffusion models (e.g., HunyuanVideo [28], Wan [48], etc.) can be finetuned with FramePack. Our experiments reveal several findings: because next-frame prediction generates smaller tensor sizes per step compared to full-video generation, it enables more balanced diffusion schedulers with less extreme flow shift timesteps. We also show that the efficient implementations of FramePack can process thousands of frames with 13B models even on laptops (e.g., 6GB or 8GB GPU memory).
@@ 2.1 Anti-forgetting and Anti-drifting
The trade-off between forgetting and drifting is also evidenced by previous discussions. CausVid [65] shows that when the video generator is causal, the quality degradation appears at the end of the video and the high-quality part may be subject to an upper bound length. DiffusionForcing [6] discussed that the cause of this drift may be related to error accumulation in models’ observation disparity between training and inference. Wang et al[51] discussed that a model with stronger memory may suffer more from drifting and error accumulation.
Noise scheduling and augmentation in history frames modify noise levels at specific timesteps, video times, or image frequencies to mitigate drifting. These methods generally reduce the dependency on past frames. DiffusionForcing [6] and RollingDiffusion [39] are typical examples. Our ablation studies investigate the influence of adding noise to history frames.
@@
Compressing latent space can improve the efficiency of video diffusion models. FlexTok [2] adjusts token context length to achieve different levels of visual content compression. LTXVideo [17] shows that a highly compressed latent space can be used for diffusing videos efficiently. PyramidFlow [25] diffuses video latents in a pyramid and re-noises downsampled latents in that pyramid to reduce computation costs. FAR [16] proposes a multi-level causal attention structure to establish long-short-term causal context pacifying and KV caches. HiTVideo [76] uses hierarchical tokenizers to enhance the video generation with autoregressive language models.
Memory in world models often involves different modeling of long-term memory. Typical examples are 3D geometry like mesh and proxy [38, 50, 67, 66]. Training diffusion models with domain data can also bake the memory into the model, with full model training [1, 46] or low-rank methods [21]. Retrieval-based memory mechanisms like WorldMem [58] are efficient when the task prioritizes reconstructing history contents.
@@ 2.2 Long Video Generation
Extending video generation beyond short clips remains an open problem. LVDM [19] generates long videos using latent diffusion, while Phenaki [47] creates variable-length videos from sequences of text prompts. Gen-L-Video [49] applies temporal co-denoising for multi-text conditioned videos, and FreeNoise [37] extends pretrained models without additional training via noise rescheduling. NUWA-XL [62] implements a Diffusion-over-Diffusion architecture with coarse-to-fine processing, while Video-Infinity [44] overcomes computational constraints through distributed generation. StreamingT2V [20] produces consistent, dynamic, and extendable videos without hard cuts, and CausVid [65] transforms bidirectional models into fast autoregressive models through distillation. Recent advances include GPT-like architecture (ViD-GPT [15]), multi-event generation (MEVG [36]), attention control for multi-prompt generation (DiTCtrl [5]), precise temporal control (MinT [55]), history-based guidance (HistoryGuidance [41]), unified next-token and full-sequence diffusion (DiffusionForcing [6]), SpectralBlend temporal attention (FreeLong [33]), video autoregressive modeling (FAR [16]), and test-time training (TTT [9]). Harvey et al[18] proposes a flexible approach for modeling long contexts with dilatation (Hierarchy-2) and propagation. Generating longer videos often requires efficient architectures, e.g., linear attention [4, 59, 52, 8, 68, 26], sparse attention [56, 71, 72, 57], low-bit computation [30, 74, 29], low-bit attention [70, 69], hidden state caching [35, 31], distillation [42, 34, 63, 64], etc.
3 Packing Frame Context
@@
We consider a video generation model that predicts next frames repeatedly to form a video. For simplicity, we consider next-frame-section prediction models using Diffusion Transformers (DiTs) that generate a section $\bm{X}$ of $S$ unknown frames so that $\bm{X}\in\mathbb{R}^{S\times h\times w\times c}$, conditioned on a section $\bm{F}$ of $T$ input frames so that $\bm{F}\in\mathbb{R}^{T\times h\times w\times c}$. All definitions of frames and pixels refer to latent representations, as most modern models operate in latent space.
@@
For next-frame (or next-frame-section) prediction, $S$ is typically 1 (or a small number). We focus on the challenging case where $T\gg S$. With per-frame context length $L_{f}$ (typically $L_{f}\approx 1560$ for each 480p frame in Hunyuan/Wan/Flux), the vanilla DiT yields total context length $L=L_{f}(T+S)$. This causes a context length explosion when $T$ is large. We observe that the input frames have different importance when predicting the next frame, and we can prioritize them according to their importance.
3.1 Time Proximity Based Packing
@@
We first consider a baseline case where the temporal proximity reflects frame importance (Fig. 1-(a)). More advanced cases involving similarity-based importance are covered in §3.2. With frames temporally closer to the prediction target being more relevant, we enumerate all frames with $\bm{F}_{0}$ being the most important (e.g., the most recent) and $\bm{F}_{T-1}$ being the least (e.g., the oldest). We define a length function $\phi(\bm{F}_{i})$ that determines each frame’s context length after VAE encoding and transformer patchifying by applying progressive compression
@@
where $\lambda>1$ is a compression parameter. The frame-wise compression is achieved by manipulating the transformer’s patchify kernel size in the input layer (e.g., $\lambda=2,i=5$ means a kernel size where the product of all dims equals $2^{5}=32$ like the 3D kernel $2\times 4\times 4$, or $8\times 2\times 2$, etc.). The total context length then follows a geometric progression
@@
and when $T\to\infty$, the total context length converges to $\lim_{T\to\infty}L=(S+\frac{\lambda}{\lambda-1})$. This bounded context length makes FramePack’s compression bottleneck invariant to the input frame number $T$.
Since most hardware supports efficient matrix processing by powers of 2, we mainly discuss the case of $\lambda=2$ in this paper. Note that we can represent arbitrary compression rates by duplicating (or dropping) several specific terms in the power-of-2 sequence: considering the accumulation $\sum_{i=0}^{+\infty}\frac{1}{2^{i}}=\frac{2}{2-1}=2$, if we want to loosen it a bit, for example to $2.625$, we can duplicate the terms $\frac{1}{2}$ and $\frac{1}{8}$ so that $\frac{1}{1}+\frac{1}{2}+(\frac{1}{2})+\frac{1}{4}+\frac{1}{8}+(\frac{1}{8})+\frac{1}{16}+...=\frac{1}{2}+\frac{1}{8}+\sum_{i=0}^{+\infty}\frac{1}{2^{i}}=2.625$. Following this, one can cover arbitrary rates by converting the rate value to binary bits and then translating every bit.
Packing schedules
Figure 1: FramePack variants. We present frame packing methods using time proximity or feature similarity. We discuss several typical kernel structures. This list does not necessarily cover all popular variants, and more structures can be developed in a similar way.
The patchifying operations in most DiTs are 3D, and we denote the 3D kernel as $(p_{f},p_{h},p_{w})$ representing the steps in frame number, height, and width. A same compression rate can be achieved by multiple possible kernel sizes, e.g., the compression rate of $64$ can be achieved by $(1,8,8)$, $(4,4,4)$, $(16,2,2)$, $(64,1,1)$, etc. Compression levels can be duplicated and combined with higher compression rates. We discuss more packing ways in Fig. 1-(c): duplication allows for same kernel sizes in frame width and height, making the compression more compact; temporal kernel can compress contiguous frames into a single tensor; symmetric progression treats both beginning and ending frames as equally important.
@@
Empirical evidence shows that using independent parameters for the different input projections at multiple compression rates facilitates stabilized learning. We assign the most commonly used input compression kernels as independent neural network layers: $(2,4,4)$, $(4,8,8)$, and $(8,16,16)$. For higher compressions (e.g., at $(16,32,32)$), we first downsample (e.g., with $2\times 2\times 2$) and then use the largest kernel $(8,16,16)$. We initialize their separated weights by interpolating from the pretrained patchifying projection (e.g., the $(2,4,4)$ projection of HunyuanVideo/Wan).
@@
While in theory FramePack can process videos of arbitrary length with a fixed, invariant context length, frames may fall below a minimum unit size (e.g., a single latent pixel) when the input length becomes extremely large. We discuss 3 options to process the tail frames: (1) simply delete the tail; (2) allow each tail frame to increase the context length by a single latent pixel; (3) apply global average pooling to all tail frames and process them with the last kernel. In our tests, the visual differences between these options are relatively negligible.
@@
When encoding inputs with different compression kernels, the different context lengths require RoPE (Rotary Position Embedding) [43] alignment. RoPE generates complex numbers with real and imaginary parts for each token position across all channels, which we refer to as “phase”. We directly downsample (using average pooling) such phases to match the compression kernels.
3.2 Feature Similarity Based Packing and Hybrid Approach
The aforementioned frame ordering $\bm{F}_{0...T-1}$ can be seen as a result of sorting all history frames using their time positions. We note that such sorting can also use other metrics like feature similarity (Fig. 1-(b)). For instance, consider a typical cosine similarity
$\operatorname{sim}_{\cos}(\bm{F}_{i},\hat{\bm{X}})=\sum_{p}\frac{(\bm{F}_{i})_{p}\cdot\hat{\bm{X}}_{p}^{\top}}{\|(\bm{F}_{i})_{p}\|\|\hat{\bm{X}}_{p}\|}\,$, (3)
where the sum is taken over pixels $p$. This measures the cosine similarity between each history frame and the estimated next frame section. Sorting the history frames using $\operatorname{sim}_{\cos}(\cdot)$ will produce a permutation $\bm{F}_{0...T-1}$ with $\bm{F}_{0}$ being the most similar and $\bm{F}_{T-1}$ being the least. This permutation can directly replace the aforementioned time proximity. Since similarity-based permutation may change abruptly when processing contiguous frames, we also consider a smooth time proximity modeling
$\operatorname{sim}_{time}(\bm{F}_{i},\hat{\bm{X}})=e^{-(\operatorname{time}(\bm{F}_{i})-\operatorname{time}(\hat{\bm{X}}))^{2}}\,$, (4)
where $\operatorname{time}(\cdot)$ gets the frames’ starting time measured in seconds. Consider the weighting
[…]
where $\lambda_{time}$ is a weighting parameter. Sorting the history frames using $\operatorname{sim}_{\text{hybrid}}(\cdot)$ will produce a permutation that transits relatively smoothly as the generating window moves forward in time. This hybrid approach is suitable for world model datasets (mainly video games) that require returning to previously visited views of scenes or events. Similar sorting can also be applied to facial identity metrics to facilitate movie generation applications that emphasize consistent human actors. This method will be evaluated in ablation experiments in the supplementary materials.
4 Drift Prevention
Drifting is a common problem in next-frame prediction models where visual quality degrades as video length increases. We discuss anti-drifting approaches by adjusting the sampling processes and history representations as shown in Fig. 2.